# Resilient Distributed Optimization for Multi-Agent Cyberphysical Systems

### v2

@@ 5.1 Convergence to optimal point almost surely
Let us consider the data values $(x_{i}(t))_{i\in\mathcal{L}}$ of the dynamic (3.1), assuming that $T_{f}$ is finite. Subsequently, for every $t\geq\max\{T_{f},T_{0}\}$ all the legitimate agents participate in the dynamic (3.1) and all the malicious agents are excluded from it. Thus from time $\max\{T_{f},T_{0}\}$ the dynamic (3.1) can be captured by a dynamic of the form (5), where $t$ is is replaced with $t-\max\{T_{f},T_{0}\}$.
@@
By Corollary 1 there exists a finite time $T_{f}$ such that every legitimate agent $i$ classifies correctly all of its legitimate and malicious neighbors at all times $t\geq T_{f}$ a.s. Thus, the dynamic (3.1) including malicious agents is equivalent to the nominal dynamic (5) with the initial inputs $z_{i}(0)=x_{i}(\max\{T_{f},T_{0}\})$ where $i\in\mathcal{L}$. By Theorem 1, the dynamic (5) converges to $x_{\mathcal{L}}^{\star}$. Additionally, by Assumption 3, $x_{\mathcal{L}}(T_{f})$ is finite for every finite $T_{f}$. Applying Corollary 1 with the a.s. finiteness of $T_{f}$ concludes the proof. ∎
@@ 5.2 Convergence to optimal point in mean
By Assumption 3 and the triangle inequality, we have $\|x-y\|^{r}\leq(2\eta)^{r}<\infty$ for every $x,y\in\mathcal{X}$. Recalling Theorem 2 we can apply the Dominated Convergence Theorem (see [54, Theorem 1.6.7]) to each sequence $\{\|x_{i}(t)-x_{\mathcal{L}}^{\star}\|^{r}\}$, $i\in\mathcal{L}$, to conclude the result. ∎
6 Finite Time Analysis:
Expected Convergence Rate
@@ 6 Finite Time Analysis:
6.1 The Expected convergence rate via the correct
classification time
Utilizing Theorem 1 and Lemma 2 we can upper bound the expected suboptimality gap by upper bounding the contributions to the suboptimality gap by the event $\{T_{f}\leq m\}$ and its complement. For each $m$ this leads to an upper bound that coincides with (16) subject to a time shift with $m$ and an additive error term of the probability that $\{T_{f}>m\}$. Since for every $m$ either $\{T_{f}\leq m\}$ is correct or its complement, we can take the minimum over the derived upper bounds.
@@ 6.1 The Expected convergence rate via the correct
If Assumptions 1-5 hold, then for every $t\geq T_{0}$ we have that the average mean squared error is bounded by a decaying function such that:
@@
$\min_{m\in[T_{0}:t-1]}\left\{\min\left\{4\eta^{2},\frac{4\overline{h}\left(t-m\right)}{\mu(t-m)(t-m+1)}+4\eta^{2}p_{e}(m)\right\}\right\}.$ (17)
@@
For every $t\geq\max\{T_{0},1\}+1$ the mean squared error $\frac{1}{|\mathcal{L}|}\sum_{i\in\mathcal{L}}\mathbf{E}[\|x_{i}(t)-x^{\star}\|^{2}]$ is upper bounded by $U_{1}(t,T_{0})$ and $U_{2}(t,T_{0})$
$U_{1}(t,T_{0})\triangleq\min\left\{4\eta^{2},\frac{4\overline{h}\left(t-T_{0}\right)}{\mu(t-T_{0})(t-T_{0}+1)}+4\eta^{2}p_{e}\left(T_{0}\right)\right\},$ (18)
$U_{2}(t,T_{0})\triangleq\min\left\{4\eta^{2},\frac{16\overline{h}\left(\frac{t-T_{0}}{2}\right)}{\mu(t-T_{0})(t-T_{0}+2)}+4\eta^{2}p_{e}\left(\frac{t+T_{0}}{2}-1\right)\right\},$ (19)
@@
$U_{3}(t,T_{0})\triangleq\min\left\{4\eta^{2},\frac{4\overline{h}\left(t-m_{3}(t)\right)}{\mu\left(t-m_{3}(t)\right)\left(t-m_{3}(t)+1\right)}+\frac{4\eta^{2}}{t}\right\}.$ (20)
@@
Since $\frac{\overline{h}(t)}{t(t+1)}$ and $p_{e}(t)$ are monotonically decreasing for $t\geq 1$, choosing the values, $m=T_{0}$, $m=(t+T_{0})/2$, and $m=m_{3}(t)$ in (4), establishes the proof of Corollary 3. ∎
We note that since both $\frac{\overline{h}(t)}{t(t+1)}$ and $p_{e}(t)$ are monotonically decreasing with $t$ for all $t\geq 1$, the functions $\frac{\overline{h}(t-m)}{(t-m)(t-m+1)}$ and $p_{e}(m)$ have opposite monotonicity with respect to $m$. The three chosen values reflect this observation. The first, i.e., $m=T_{0}$, is aimed to minimize the term $\frac{\overline{h}(t-m)}{(t-m)(t-m+1)}$ at the expense of $p_{e}(m)$ which for this choice of $m=T_{0}$ is a constant with respect to $t$. This choice of $m$ is favorable when $T_{0}$ is close to $t$. The choice $m=m(t)=(t+T_{0})/2$ is seleceted so that both the terms $\frac{\overline{h}(t-m)}{(t-m)(t-m+1)}$ and $p_{e}(t)$ decrease as a function of $t$, at a linear and an exponential rate, respectively. The last choice, i.e., $m=m_{3}(t)$, balances the convergence rate of the two additive terms that compose the objective function of (4) and thus is best to use for large values of $t$.
Now, we proceed to prove Theorem 4 utilizing the law of total expectation with respect to the event $\{T_{f}\leq m\}$. Additionally, we provide in Section II of the supplementary material an alternative proof for Theorem 4 that considers the individual events $T_{f}\triangleq\max\{k,T_{0}\}$.
@@
$\stackrel{{\scriptstyle(c)}}{{\leq}}\min\left\{4\eta^{2},\frac{4\overline{h}\left(t-m\right)}{\mu(t-m)(t-m+1)}\right\}+4\eta^{2}p_{e}(m),$
where $(a)$ follows from the law of total expectation, $(b)$ follows from the definition of $T_{f}$ and Assumption 3, and $(c)$ follows from the fact that $\frac{\overline{h}(t)}{t(t+1)}$ is nonincreasing for $t\geq 1$ and Lemma 2. ∎
6.2 Convergence results via trustworthiness
@@
$\mathbf{E}\bigg{[}\|c_{i}(t)-\overline{c}_{i}(t)\|^{r}\bigg{]}\leq(2\eta)^{r}p_{c}(t)$
@@
$+\Pr(W_{\mathcal{L}}(t)\neq\overline{W}_{\mathcal{L}})\mathbf{E}\bigg{[}\|c_{i}(t)-\overline{c}_{i}(t)\|^{r}\big{|}W_{\mathcal{L}}(t)\neq\overline{W}_{\mathcal{L}}\bigg{]}$
$\leq(2\eta)^{r}p_{c}(t).$
@@
Let $\mathcal{X}\subset\mathbb{R}^{d}$ be a compact and convex set. Additionally, let $W(t)\triangleq(w_{ij}(t))_{i,j\in\mathcal{L}}$ be deterministic doubly stochastic matrices such that $\sigma_{2}(W(t))\leq\rho$ for all $t\geq 0$. Furthermore, let $\Delta_{i}(t)\in\mathbb{R}^{d\times 1}$ be random vectors, and let $X(t)\in\mathbb{R}^{d\times|\mathcal{L}|}$ be defined by the following dynamic
@@
Then,
@@
$\hskip 2.84544pt+\frac{2(\eta\sqrt{p_{c}((t+T_{0})/2)}+G\gamma((t-T_{0})/2))}{1-\rho_{\mathcal{L}}}.$ (22)
@@
$+2\sqrt{\mathbf{E}\left[\|c_{i}(t)-\overline{c}_{i}(t)\|^{2}\right]}\sqrt{\mathbf{E}\left[\|\phi_{i}(t)\|^{2}\right]}.$
@@
$\lim_{t\rightarrow\infty}\mathbf{E}\left[\|x_{i}(t)-x_{\mathcal{L}}^{\star}\|^{2}\right]=0,\>\forall\>i\in\mathcal{L}.$ (24)
Moreover, let $\gamma(t)=\frac{2}{\mu(t+2)}$. Then, for every $T_{0}\geq 0$ and $T_{0}$ there exists a function $C_{\mathcal{M}}(T_{0})$ that decreases exponentially with $T_{0}$ and is independent of $T$ such that for any collection $x_{i}(0)\in\mathcal{X}$, $i\in\mathcal{L}$, and for all $T\geq T_{0}$,
$\frac{1}{|\mathcal{L}|}\sum_{i\in\mathcal{L}}\mathbf{E}\left[\|x_{i}(T)-x_{\mathcal{L}}^{\star}\|^{2}\right]\leq\min\left\{4\eta^{2},\frac{4\overline{h}(T-T_{0})+C_{\mathcal{M}}(T_{0})}{\mu(T-T_{0})(T-T_{0}+1)}\right\},$ (25)
where $\overline{h}(\cdot)$ is defined in (5).
@@
$4\eta^{2}p_{c}(t)\left[2(L+1)+\gamma^{2}(t-T_{0})L^{2}+\frac{\gamma(t-T_{0})G(L+1)}{2\eta}\right],$
and recall that $\overline{x}_{\mathcal{L}}(t)\triangleq\frac{1}{|\mathcal{L}|}\sum_{i\in\mathcal{L}}x_{i}(t)$. Therefore,
@@
$\leq h_{\mathcal{M}}(t,T_{0})+\frac{\left(1-\mu\gamma(t-T_{0})\right)}{|\mathcal{L}|}\sum_{i\in\mathcal{L}}\mathbf{E}\left[\|x_{i}(t)-x_{\mathcal{L}}^{\star}\|^{2}\right]$
@@
$\hskip 99.58464pt+\frac{\mu+L}{2}\mathbf{E}\left[\|x_{i}(t)-\overline{x}_{\mathcal{L}}(t)\|^{2}\right]\Bigg{]},$
where the last inequality follows from Assumption 4, the convexity of $\|\cdot\|^{2}$ and the double stochasticity of $\overline{W}_{\mathcal{L}}$.
Recall (22) and denote
$\widetilde{h}_{\mathcal{M}}(t,T_{0})\triangleq\gamma(t-T_{0})G^{2}+2G\delta_{\mathcal{M}}(t,T_{0})+(\mu+L)\delta_{\mathcal{M}}^{2}(t,T_{0}).$
@@
$+h_{\mathcal{M}}(t,T_{0})+\frac{\left(1-\mu\gamma(t-T_{0})\right)}{|\mathcal{L}|}\sum_{i\in\mathcal{L}}\mathbf{E}\left[\|x_{i}(t)-x_{\mathcal{L}}^{\star}\|^{2}\right].$
@@
To prove the second part of the theorem, we let $\gamma(t)=\frac{2}{\mu(t+2)}$ as proposed in [58]. It follows that
$\frac{1}{|\mathcal{L}|}\sum_{i\in\mathcal{L}}\mathbf{E}\left[\|x_{i}(t+1)-x_{\mathcal{L}}^{\star}\|^{2}\right]\leq\frac{2\widetilde{h}_{\mathcal{M}}(t,T_{0})}{\mu(t-T_{0}+2)}$
$\>\>+h_{\mathcal{M}}(t,T_{0})+\frac{t}{t-T_{0}+2}\cdot\frac{1}{|\mathcal{L}|}\sum_{i\in\mathcal{L}}\mathbf{E}\left[\|x_{i}(t)-x_{\mathcal{L}}^{\star}\|^{2}\right].$
@@
$\frac{1}{|\mathcal{L}|}\sum_{i\in\mathcal{L}}\mathbf{E}[\|x_{i}(T)-x_{\mathcal{L}}^{\star}\|^{2}]\leq\frac{2\sum_{t=T_{0}}^{T-1}(t-T_{0}+1)\widetilde{h}_{\mathcal{M}}(t,T_{0})}{\mu(T-T_{0})(T-T_{0}+1)}$
@@
In addition, we utilize the identities for $|v|\in(0,1)$
$\sum_{t=0}^{\infty}tv^{t}=\frac{v}{(1-v)^{2}}\text{ and }\sum_{t=0}^{\infty}t^{2}v^{t}=\frac{2v}{(1-v)^{3}}.$
Denote
$\widetilde{C}_{1}(T_{0},E,D)\triangleq\frac{16\eta e^{-T_{0}E^{2}}\sqrt{D}}{1-\rho_{\mathcal{L}}}\Bigg{[}\frac{G}{(1-e^{-E^{2}})^{2}}+$
$\frac{G+(\eta+\frac{4}{\mu})(\mu+L)}{(1-\rho_{\mathcal{L}})^{2}}+\frac{(\mu+L)(G+2\mu\sqrt{D}e^{-T_{0}E^{2}})}{\mu(1-\rho_{\mathcal{L}})^{3}}\Bigg{]},$
@@
$\sum_{t=T_{0}}^{T-1}(t-T_{0}+1)\widetilde{h}_{\mathcal{M}}(t,T_{0})\leq 2\overline{h}(t-T_{0})+C_{1}(T_{0}),$
@@
$\widetilde{C}_{2}(T_{0},E,D)\triangleq\frac{4\eta(L+1)De^{-2T_{0}E^{2}}}{\left(1-e^{-2E^{2}}\right)}$
$\>\>\cdot\Bigg{[}\frac{4\eta e^{-2E^{2}}}{\left(1-e^{-2E^{2}}\right)^{2}}+\frac{(6\eta+\frac{G}{\mu})e^{-2E^{2}}}{1-e^{-2E^{2}}}+4\eta+\frac{4\eta L}{\mu^{2}}+\frac{G}{\mu}\Bigg{]},$
and $C_{2}(T_{0})\triangleq\widetilde{C}_{2}(T_{0},E_{\mathcal{L}},D_{\mathcal{L}})+\widetilde{C}_{2}(T_{0},E_{\mathcal{M}},D_{\mathcal{M}})$.
Then,
$\sum_{t=T_{0}}^{T-1}(t-T_{0}+1)(t-T_{0}+2)h_{\mathcal{M}}(t,T_{0})$
$=8\eta^{2}(L+1)\sum_{t=0}^{T-T_{0}-1}(t+1)(t+2)p_{c}(t+T_{0})$
$+4\eta^{2}L^{2}\sum_{t=0}^{T-T_{0}-1}(t+1)(t+2)\gamma^{2}(t)p_{c}(t+T_{0})$
$+2\eta G(L+1)\sum_{t=0}^{T-T_{0}-1}(t+1)(t+2)\gamma(t)p_{c}(t+T_{0})\leq C_{2}(T_{0}).$

### v3

@@ 5.1 Convergence to optimal point almost surely
Let us consider the data values $(x_{i}(t))_{i\in\mathcal{L}}$ of the dynamic (3.1), assuming that $T_{f}$ is finite. Subsequently, for every $t\geq\max\{T_{f},T_{0}\}$ all the legitimate agents participate in the dynamic (3.1) and all the malicious agents are excluded from it. Thus, from time $\max\{T_{f},T_{0}\}$ the dynamic (3.1) can be captured by a dynamic of the form (5), where $t$ is replaced with $t-\max\{T_{f},T_{0}\}$.
@@
By Corollary 1 there exists a finite time $T_{f}$ such that every legitimate agent $i$ classifies correctly all of its legitimate and malicious neighbors at all times $t\geq T_{f}$ a.s. Thus, during this time interval, the dynamic (3.1) is equivalent to the nominal dynamic (5) with the initial inputs $z_{i}(0)=x_{i}(\max\{T_{f},T_{0}\})$ where $i\in\mathcal{L}$. By Theorem 1, the dynamic (5) converges to $x_{\mathcal{L}}^{\star}$. Additionally, by Assumption 3, $x_{\mathcal{L}}(T_{f})$ is finite for every finite $T_{f}$. Applying Corollary 1 with the a.s. finiteness of $T_{f}$ concludes the proof. ∎
@@ 5.2 Convergence to optimal point in mean
By the triangle inequality and Assumption 3, we have $\|x-y\|^{r}\leq(2\eta)^{r}<\infty$ for all $x,y\in\mathcal{X}$. Recalling Theorem 2 we can apply the Dominated Convergence Theorem (see [54, Theorem 1.6.7]) to each sequence $\{\|x_{i}(t)-x_{\mathcal{L}}^{\star}\|^{r}\}$, $i\in\mathcal{L}$, to conclude the result. ∎
6 Finite Time Analysis: Expected Convergence Rate
@@
6.1 Convergence rate via correct classification time
Utilizing Theorem 1 and Lemma 2 we can upper bound the expected suboptimality gap by upper bounding the contributions to the suboptimality gap by the event $\{T_{f}\leq m\}$ and its complement. For each $m$ this leads to an upper bound that coincides with (16) subject to a time shift with $m$ and an additive error term of the probability that $\{T_{f}>m\}$. Since for every $m$ either $\{T_{f}\leq m\}$ is correct or its complement, we can take the minimum over the derived upper bounds. Before presenting the resulting theorem, recall that $\mu_{\epsilon}\triangleq\mu(1-\epsilon)$ with $\epsilon\in(0,1)$, and that $\eta$ is the maximal input norm, i.e. $\|x_{i}(t)\|\leq\eta$ for every $i\in\mathcal{L}\cup\mathcal{M}$. Additionally, recall that $\overline{h}_{\epsilon}(\cdot)$ is defined in (5) and dictates the convergence rate of the nominal dynamic (5). Finally recall the misclassification probability bound $p_{e}(k)$ defined in (10).
@@ 6.1 Convergence rate via correct classification time
Let $\gamma(t)=\frac{2}{\mu_{\epsilon}(t+2)}$. If Assumptions 1-5 hold, then for every $t\geq T_{0}$ we have that the average mean squared error is bounded by a decaying function such that for any $\epsilon\in(0,1)$:
@@
$\min_{m\in[T_{0}:t-1]}\left\{\min\left\{4\eta^{2},\frac{4\overline{h}_{\epsilon}\left(t-m\right)}{\mu_{\epsilon}(t-m)(t-m+1)}+4\eta^{2}p_{e}(m)\right\}\right\}.$ (17)
@@
Let $\gamma(t)=\frac{2}{\mu_{\epsilon}(t+2)}$. For every $t\geq\max\{T_{0},1\}+1$ and $\epsilon\in(0,1)$ the mean squared error $\frac{1}{|\mathcal{L}|}\sum_{i\in\mathcal{L}}\mathbf{E}[\|x_{i}(t)-x^{\star}\|^{2}]$ is upper bounded by $U_{1}(t,T_{0})$ and $U_{2}(t,T_{0})$
$U_{1}(t,T_{0})\triangleq\min\left\{4\eta^{2},\frac{4\overline{h}_{\epsilon}\left(t-T_{0}\right)}{\mu_{\epsilon}(t-T_{0})(t-T_{0}+1)}+4\eta^{2}p_{e}\left(T_{0}\right)\right\},$ (18)
$U_{2}(t,T_{0})\triangleq\min\left\{4\eta^{2},\frac{16\overline{h}_{\epsilon}\left(\frac{t-T_{0}}{2}\right)}{\mu_{\epsilon}(t-T_{0})(t-T_{0}+2)}+4\eta^{2}p_{e}\left(\frac{t+T_{0}}{2}-1\right)\right\},$ (19)
@@
$U_{3}(t,T_{0})\triangleq\min\left\{4\eta^{2},\frac{4\overline{h}_{\epsilon}\left(t-m_{3}(t)\right)}{\mu_{\epsilon}\left(t-m_{3}(t)\right)\left(t-m_{3}(t)+1\right)}+\frac{4\eta^{2}}{t}\right\},$ (20)
for any $\epsilon\in(0,1)$.
@@
Since $\frac{\overline{h}_{\epsilon}(t)}{t(t+1)}$ and $p_{e}(t)$ are monotonically decreasing for $t\geq 1$, choosing the values, $m=T_{0}$, $m=(t+T_{0})/2$, and $m=m_{3}(t)$ in (4), establishes the proof of Corollary 3. ∎
We note that since both $\frac{\overline{h}_{\epsilon}(t)}{t(t+1)}$ and $p_{e}(t)$ are monotonically decreasing with $t$ for all $t\geq 1$, and $\mu_{\epsilon}>0$ the functions $\frac{\overline{h}_{\epsilon}(t-m)}{(t-m)(t-m+1)}$ and $p_{e}(m)$ have opposite monotonicity with respect to $m$. The three chosen values reflect this observation. The first, i.e., $m=T_{0}$, is aimed to minimize the term $\frac{\overline{h}_{\epsilon}(t-m)}{(t-m)(t-m+1)}$ at the expense of $p_{e}(m)$ which for this choice of $m=T_{0}$ is a constant with respect to $t$. This choice of $m$ is favorable when $T_{0}$ is close to $t$. The choice $m=m(t)=(t+T_{0})/2$ is selected so that both the terms $\frac{\overline{h}_{\epsilon}(t-m)}{(t-m)(t-m+1)}$ and $p_{e}(t)$ decrease as a function of $t$, at a linear and an exponential rate, respectively. The last choice, i.e., $m=m_{3}(t)$, balances the convergence rate of the two additive terms that compose the objective function of (4) and thus is best to use for large values of $t$.
Now, we proceed to prove Theorem 4 utilizing the law of total expectation with respect to the event $\{T_{f}\leq m\}$. Additionally, we provide in Section II of the supplementary material an alternative proof for Theorem 4 that considers the individual events $M_{k}\triangleq\max\{k,T_{0}\}$.
@@
$\stackrel{{\scriptstyle(c)}}{{\leq}}\min\left\{4\eta^{2},\frac{4\overline{h}_{\epsilon}\left(t-m\right)}{\mu_{\epsilon}(t-m)(t-m+1)}\right\}+4\eta^{2}p_{e}(m),$
where $(a)$ follows from the law of total expectation, $(b)$ follows from the definition of $T_{f}$ and Assumption 3, and $(c)$ follows from the fact that $\frac{\overline{h}_{\epsilon}(t)}{t(t+1)}$ is nonincreasing for $t\geq 1$ and Lemma 2. ∎
6.2 Convergence rate via trustworthiness
@@
$\mathbf{E}\bigg{[}\|c_{i}(t)-\overline{c}_{i}(t)\|^{r}\bigg{]}\leq(2\eta)^{r}p_{c}(t).$
@@
$\quad+\Pr(W_{\mathcal{L}}(t)\neq\overline{W}_{\mathcal{L}})\mathbf{E}\bigg{[}\|c_{i}(t)-\overline{c}_{i}(t)\|^{r}\big{|}W_{\mathcal{L}}(t)\neq\overline{W}_{\mathcal{L}}\bigg{]}$
$\leq(2\eta)^{r}p_{c}(t).\hskip 170.71652pt\qed$
@@
Let $\mathcal{X}\subset\mathbb{R}^{d}$ fulfill Assumption 3. Additionally, let $W(t)\triangleq(w_{ij}(t))_{i,j\in\mathcal{L}}$ be deterministic doubly stochastic matrices such that $\sigma_{2}(W(t))\leq\rho$ for all $t\geq 0$. Furthermore, let $\Delta_{i}(t)\in\mathbb{R}^{d\times 1}$ be random vectors, and let $X(t)\in\mathbb{R}^{d\times|\mathcal{L}|}$ be defined by the following dynamic
@@
Then, for every $X(0)$ such that $[X(0)]_{i}\in\mathcal{X},\>\>\forall\>i\in\mathcal{L}$:
@@
$\quad+\frac{2(\eta\sqrt{p_{c}((t+T_{0})/2)}+G\gamma((t-T_{0})/2))}{1-\rho_{\mathcal{L}}}.$ (22)
@@
$\quad+2\sqrt{\mathbf{E}\left[\|c_{i}(t)-\overline{c}_{i}(t)\|^{2}\right]}\sqrt{\mathbf{E}\left[\|\phi_{i}(t)\|^{2}\right]}.$
@@
$\lim_{t\rightarrow\infty}\mathbf{E}\left[\|x_{i}(t)-x_{\mathcal{L}}^{\star}\|^{2}\right]=0,\quad\forall i\in\mathcal{L}.$ (24)
Moreover, let $\gamma(t)=\frac{2}{\mu_{\epsilon}(t+2)}$. Then, for every $T_{0}\geq 0$ and $\epsilon\in(0,1)$ there exists a function $C_{\mathcal{M}}(T_{0})$ that decreases exponentially with $T_{0}$ and is independent of $T$ such that for any collection $x_{i}(0)\in\mathcal{X}$, $i\in\mathcal{L}$, $\epsilon\in(0,1)$ and for all $T\geq T_{0}$,
$\frac{1}{|\mathcal{L}|}\sum_{i\in\mathcal{L}}\mathbf{E}\left[\|x_{i}(T)-x_{\mathcal{L}}^{\star}\|^{2}\right]\leq\min\left\{4\eta^{2},\frac{4\overline{h}_{\epsilon}(T-T_{0})+C_{\mathcal{M}}(T_{0})}{\mu_{\epsilon}(T-T_{0})(T-T_{0}+1)}\right\},$ (25)
where $\overline{h}_{\epsilon}(\cdot)$ is defined in (5) and $\mu_{\epsilon}\triangleq\mu(1-\epsilon)$.
@@
$h_{\mathcal{M}}(t,T_{0})\triangleq 4\eta^{2}p_{c}(t)\left[3+\frac{(G+4L\eta)\gamma(t-T_{0})}{\eta}+\frac{L(G+L\eta)\gamma^{2}(t-T_{0})}{\eta}\right],$
and recall that $\overline{x}_{\mathcal{L}}(t)\triangleq\frac{1}{|\mathcal{L}|}\sum_{i\in\mathcal{L}}x_{i}(t)$. Therefore, let $\epsilon\in(0,1)$
@@
$\leq h_{\mathcal{M}}(t,T_{0})+\frac{\left(1-\mu(1-\epsilon)\gamma(t-T_{0})\right)}{|\mathcal{L}|}\sum_{i\in\mathcal{L}}\mathbf{E}\left[\|x_{i}(t)-x_{\mathcal{L}}^{\star}\|^{2}\right]$
@@
$\hskip 99.58464pt+\frac{\mu(1/\epsilon-1)+L}{2}\mathbf{E}\left[\|x_{i}(t)-\overline{x}_{\mathcal{L}}(t)\|^{2}\right]\Bigg{]},$
where the last inequality follows from Assumption 4, the convexity of $\|\cdot\|^{2}$, the double stochasticity of $\overline{W}_{\mathcal{L}}$, and Young inequality (for further details see the proof of Theorem 1).
Recall (22) and the notation $\mu_{\epsilon}\triangleq\mu(1-\epsilon)$. Denote
$\widetilde{h}_{\mathcal{M}}(t,T_{0})\triangleq\gamma(t-T_{0})G^{2}+2G\delta_{\mathcal{M}}(t,T_{0})+(\mu_{\epsilon}/\epsilon+L)\delta_{\mathcal{M}}^{2}(t,T_{0}).$
@@
$+h_{\mathcal{M}}(t,T_{0})+\frac{\left(1-\mu_{\epsilon}\gamma(t-T_{0})\right)}{|\mathcal{L}|}\sum_{i\in\mathcal{L}}\mathbf{E}\left[\|x_{i}(t)-x_{\mathcal{L}}^{\star}\|^{2}\right].$
@@
To prove the second part of the theorem, motivated by [58] we let $\gamma(t)=\frac{2}{\mu_{\epsilon}(t+2)}$. It follows that
$\frac{1}{|\mathcal{L}|}\sum_{i\in\mathcal{L}}\mathbf{E}\left[\|x_{i}(t+1)-x_{\mathcal{L}}^{\star}\|^{2}\right]\leq\frac{2\widetilde{h}_{\mathcal{M}}(t,T_{0})}{\mu_{\epsilon}(t-T_{0}+2)}$
$\>\>+h_{\mathcal{M}}(t,T_{0})+\frac{t-T_{0}}{t-T_{0}+2}\cdot\frac{1}{|\mathcal{L}|}\sum_{i\in\mathcal{L}}\mathbf{E}\left[\|x_{i}(t)-x_{\mathcal{L}}^{\star}\|^{2}\right].$
@@
$\frac{1}{|\mathcal{L}|}\sum_{i\in\mathcal{L}}\mathbf{E}[\|x_{i}(T)-x_{\mathcal{L}}^{\star}\|^{2}]\leq\frac{2\sum_{t=T_{0}}^{T-1}(t-T_{0}+1)\widetilde{h}_{\mathcal{M}}(t,T_{0})}{\mu_{\epsilon}(T-T_{0})(T-T_{0}+1)}$
@@
In addition, we utilize the identities for $v\in(0,1)$
$\sum_{t=0}^{\infty}tv^{t}=\frac{v}{(1-v)^{2}}\text{ and }\sum_{t=0}^{\infty}t^{2}v^{t}=\frac{v(v+1)}{(1-v)^{3}}.$
Denote,
$\widetilde{C}_{1}(T_{0},E,D)\triangleq\frac{16\eta e^{-T_{0}E^{2}}\sqrt{D}}{1-\rho_{\mathcal{L}}}\Bigg{[}\frac{2\eta(1-\rho_{\mathcal{L}})+G/\epsilon+L/\mu_{\epsilon}}{(1-\rho_{\mathcal{L}})(1-\rho_{\mathcal{L}}e^{-E^{2}})}+$
$\frac{G+2(\eta+\frac{G}{\mu_{\epsilon}})(\mu_{\epsilon}/\epsilon+L)}{(1-\rho_{\mathcal{L}})^{2}}+\frac{(\mu_{\epsilon}/\epsilon+L)(G/\mu_{\epsilon}+2\eta\sqrt{D}e^{-T_{0}E^{2}})}{(1-\rho_{\mathcal{L}})^{3}}$
$\frac{G}{(1-e^{-E^{2}})^{2}}+\frac{2(\mu_{\epsilon/\epsilon+L})(G(1-e^{-E^{2}})+\eta\sqrt{D}e^{-T_{0}E^{2}})}{(1-\rho_{\mathcal{L}})(1-e^{-E^{2}})^{2}}\Bigg{]},$
@@
$\sum_{t=T_{0}}^{T-1}(t-T_{0}+1)\widetilde{h}_{\mathcal{M}}(t,T_{0})\leq 2\overline{h}_{\epsilon}(t-T_{0})+C_{1}(T_{0}),$
@@
$\widetilde{C}_{2}(T_{0},E,D)\triangleq\frac{8\eta De^{-2T_{0}E^{2}}}{\left(1-e^{-2E^{2}}\right)}$
$\>\>\cdot\Bigg{[}\frac{3\eta}{\left(1-e^{-2E^{2}}\right)^{2}}+\frac{G+4L\eta}{\mu_{\epsilon}(1-e^{-2E^{2}})}+\frac{2L(G+L\eta)}{\mu_{\epsilon}^{2}}\Bigg{]},$
and $C_{2}(T_{0})\triangleq\widetilde{C}_{2}(T_{0},E_{\mathcal{L}},D_{\mathcal{L}})+\widetilde{C}_{2}(T_{0},E_{\mathcal{M}},D_{\mathcal{M}})$. Then,
$\sum_{t=T_{0}}^{T-1}(t-T_{0}+1)(t-T_{0}+2)h_{\mathcal{M}}(t,T_{0})\leq C_{2}(T_{0}).$